OpenStudy (dancergirl45):

Write the equation of the line that passes through the points (3, 6) and (4, 10) using function notation. f(x)=4x-6 f(x)=x+4 y=x+4 y=4x-6

OpenStudy (dancergirl45):

@graysongraddylol @SageWilson

OpenStudy (dancergirl45):

sage please help

Vocaloid (vocaloid):

start with the slope formula

Vocaloid (vocaloid):

m = (y2-y1)/(x2-x1) = ?

OpenStudy (dancergirl45):

4-3 10-6

Vocaloid (vocaloid):

keep going (10-6)/(4-3) = ?

OpenStudy (loginusernamed):

5 1/8 + 5/6 + 5/12

OpenStudy (dancergirl45):

4 1

Vocaloid (vocaloid):

keep going... 4/1 = ?

OpenStudy (dancergirl45):

whattt

Vocaloid (vocaloid):

/ means divide....

OpenStudy (dancergirl45):

4

Vocaloid (vocaloid):

good, so that's our slope now we plug in the slope into the equation y = mx + b and solve for b 6 = 4(3) + b b = ?

OpenStudy (dancergirl45):

12

Vocaloid (vocaloid):

nope.

Vocaloid (vocaloid):

6 = 12 + b b = ?

OpenStudy (dancergirl45):

2

Vocaloid (vocaloid):

nope, please check your math...

Vocaloid (vocaloid):

12 + ___ = 6 fill in the blank

OpenStudy (dancergirl45):

-6

Vocaloid (vocaloid):

good, so b = -6 and m = 4 so our equation is y = 4x - 6

OpenStudy (dancergirl45):

Thank you so much!!

OpenStudy (dancergirl45):

This was wrong

Vocaloid (vocaloid):

oh whoops, they wanted function notation so f(x) =4x - 6

OpenStudy (dancergirl45):

thanks after i submitted it